OpenStudy (anonymous):

Two pipes empty into a tank. One pipe can fill the tank in 10.0 h, and the other in 11.0 h. How long will it take both pipes together to fill the tank? Round your answer to 1 decimal place. Both pipes together will fill the tank in ? h

OpenStudy (anonymous):

\[\frac{10\times 11}{10+11}\]

OpenStudy (anonymous):

that is the quick way i can explain the long way if you like, but that will be what we will get

OpenStudy (anonymous):

so how would you figure out the answer from this

OpenStudy (anonymous):

you mean how did i arrive at what i wrote, or how do you compute it?

OpenStudy (anonymous):

i get \[\frac{110}{21}\] or as a decimal \(5.2\) rounded

OpenStudy (anonymous):

i can derive that answer for you if that is what you are asking

OpenStudy (anonymous):

ready?